how much water should be added to a 20L of a 20% salt solution to dilute it to 15% salt solution
Answer by stanbon(75887)

how much water should be added to a 20L of a 20% salt solution to dilute it to 15% salt solution
------
Equation:
water + water = water
0.80*20 + 1*x = 0.85(20+x)
----
80*20 + 100x = 85*20 + 85x
15x = 5*20
x = 6 2/3 L (amt of water to add)
--------------
Cheers,
Stan H.
---------------